what is the slope and y-intercept of:
5-3y=7x
-2x+4-3=0
9x-2-(7/3)=0
8-(x/3)-(y/5)=0

Answer by Alan3354(69443)

what is the slope and y-intercept of:
5-3y=7x
-2x+4-3=0
9x-2-(7/3)=0
8-(x/3)-(y/5)=0
----------
Put them in slope-intercept form -- y = mx + b
m = slope
b = y-intercept
-----
That might sound complicated, but all you do is solve for y.
----
5-3y=7x
y = (-7/3)x + 5
m = -7/3, b = 5
-------
-2x+4-3=0
There's no y term in this one.
x = 1/2
For all values of y, x = 1/2
That's parallel to the y-axis, or "vertical"
--> m is undefined, or infinite
Since it's parallel, it doesn't intercept the y-axis, there's on b
------------
9x-2-(7/3)=0
Same as above, no y
--------------
8-(x/3)-(y/5)=0
Solve for y, same as #1